Estimate the time required for a 5 MeV alpha particle to slow down


and stop in silicon. Repeat for the same particle in hydrogen gas?

1
Expert's answer
2022-01-05T04:21:36-0500

In this problem, we can use the equation from the textbook:


"T_\text{Si}=\dfrac{1.2R\sqrt{m/E}}{10^7},\\\space\\
T_\text{Si}=\dfrac{1.2(22·10^{-6})\sqrt{4/5}}{10^7}=2.4\text{ ns}.\\\space\\
T_\text{Si}=\dfrac{1.2R\sqrt{m/E}}{10^7},\\\space\\
T_\text{Si}=\dfrac{1.2(0.1)\sqrt{4/5}}{10^7}=11\text{ ns}.\\\space\\"